OpenStudy (oreolover4876):

hey im studying for my geometry finals that are sadly tomorrow and I cant remember how to construct or do proportions. Please help!

OpenStudy (anonymous):

I can't help you with the construct part, but I can help you with proportions!

OpenStudy (oreolover4876):

That's fine any help would be great.

OpenStudy (anonymous):

Ok, so ask me a question about proportions, and I'll show you how to solve it.

OpenStudy (anonymous):

Never mind, I'll just make up a proportion question and show you how to complete it.

OpenStudy (oreolover4876):

kay that's fine this laptop is really slow.

OpenStudy (oreolover4876):

 I have to solve and create proportions based off of these triangles

OpenStudy (anonymous):

Ok, so for the first question 9/24=3/b. The easiest way to solve problems like this is to cross multiply. So, we multiply 24 and 3. So 24*3=9*b. The answer to 24*3 is 72. We divide each side by 9. 9*b/9=b, so now what we have left is 72 divided by 9=b. 72/9=8, so b=8.

OpenStudy (anonymous):

Ok, to solve the problem you have there, we have to create the proportion first to solve it.

OpenStudy (anonymous):

The proportion would be 14/18=x/12.

OpenStudy (anonymous):

If you want to know how I found the proportion from the triangles, I used the 14 from the side of the first triangle and set it as the numerator and 18 as the denominator, and I set x as the numerator and 12 as the denominator.

OpenStudy (anonymous):

And to solve that proportion, just cross multiply like I did with the other equation. I hope this helps!

OpenStudy (oreolover4876):

sorry for the late response yes that helps a lot thank you soooo much!